# X-IONet: Cross-Platform Inertial Odometry Network with Dual-Stage Attention

### v1

@@ Abstract
Learning-based inertial odometry has achieved remarkable progress in pedestrian navigation. However, extending these methods to quadruped robots remains challenging due to their distinct and highly dynamic motion patterns. Models that perform well on pedestrian data often experience severe degradation when deployed on legged platforms. To tackle this challenge, we introduce X-IONet, a cross-platform inertial odometry framework that operates solely using a single Inertial Measurement Unit (IMU). X-IONet incorporates a rule-based expert selection module to classify motion platforms and route IMU sequences to platform-specific expert networks. The displacement prediction network features a dual-stage attention architecture that jointly models long-range temporal dependencies and inter-axis correlations, enabling accurate motion representation. It outputs both displacement and associated uncertainty, which are further fused through an Extended Kalman Filter (EKF) for robust state estimation. Extensive experiments on public pedestrian datasets and a self-collected quadruped robot dataset demonstrate that X-IONet achieves state-of-the-art performance, reducing Absolute Trajectory Error (ATE) by 14.3% and Relative Trajectory Error (RTE) by 11.4% on pedestrian data, and by 52.8% and 41.3% on quadruped robot data. These results highlight the effectiveness of X-IONet in advancing accurate and robust inertial navigation across both human and legged robot platforms.
@@ I INTRODUCTION
Inertial navigation estimates ego-motion and position solely from Inertial Measurement Unit (IMU) data. Unlike visual–inertial odometry [1, 2], lidar–inertial odometry [3, 4, 5], or GNSS-based localization [6, 7], inertial navigation is fully self-contained and does not rely on external signals [8]. This independence enables robust performance in visually degraded environments and GNSS-denied settings. However, unavoidable sensor noise causes significant drift during long-term pure-inertial integration, making accurate standalone inertial navigation a long-standing challenge.
@@
Recently, deep learning has shown great potential to enhance inertial odometry. Learning-based methods leverage neural architectures to directly extract motion patterns from IMU signals, effectively constraining dead-reckoning drift and improving state estimation accuracy. These approaches have been applied to pedestrian motion [9, 10, 11, 12, 13, 14], legged robots [15], and unmanned aerial vehicles (UAVs) [16, 17, 18, 19].
@@
Despite promising progress, most existing methods are tailored to a single platform, leaving cross-platform inertial odometry largely unsolved. Although pedestrian and quadruped locomotion share certain characteristics—such as periodicity and gravity-related priors—quadruped motion is considerably more dynamic and diverse. As shown in Fig. 1, quadrupeds perform rapid accelerations and decelerations, lateral and backward movements, and highly non-periodic maneuvers, producing complex inertial signatures absent in human motion. Moreover, inertial motion cues are inherently multi-scale and multi-dimensional; existing architectures struggle to extract and fuse such features effectively, resulting in inaccurate displacement predictions. These challenges call for new neural architectures that capture both long-range temporal dependencies and cross-axis correlations in IMU data.
Figure 1: Cross-Platform Inertial Odometry Network for pedestrians and quadruped robots. The temporal and frequency-domain visualizations reveal distinct differences between pedestrian and quadruped inertial signals.
@@
To address these challenges, we propose X-IONet (Cross-Platform Inertial Odometry Network), a novel and generalizable deep learning framework for inertial odometry estimation across both human and quadruped robotic platform. X-IONet incorporates a rule-based expert selection module that routes IMU data to specialized expert networks based on platform characteristics. Each expert network adopts a dual-stage attention–based displacement prediction model, capturing long-range temporal dependencies via temporal self-attention and inter-axis relationships via dimensional self-attention. The network predicts motion displacement and corresponding uncertainty, and integrates these estimates into an Extended Kalman Filter (EKF) for robust motion state estimation. Extensive experiments on public pedestrian datasets and a newly collected quadruped dataset show that X-IONet achieves state-of-the-art accuracy, exhibits strong cross-platform generalization, and maintains robustness under aggressive and irregular motion patterns.
@@
  • We conduct comprehensive evaluations on both pedestrian and quadruped datasets, demonstrating that X-IONet generalizes across distinct motion patterns and achieves state-of-the-art performance in inertial navigation tasks.
@@ II RELATED WORK
Recent years have witnessed rapid progress in deep learning–based inertial odometry, where data-driven neural networks are trained to directly model inertial signals for accurate motion state estimation.
@@
IONet [9] first demonstrated the feasibility of learning motion patterns directly from raw inertial data using recurrent neural networks. RoNIN [14] further advanced this direction by proposing three end-to-end frameworks based on Residual Networks (ResNet) [20], Long Short-Term Memory (LSTM) networks [21], and Temporal Convolutional Networks (TCN) [22], and also released a large-scale dataset for pedestrian inertial navigation. With the success of Transformer architectures in natural language processing [23] and computer vision [24], CTIN [12] introduced a Transformer-based model for inertial data, enabling global motion pattern modeling through dynamic attention computation. Building upon this idea, iMoT [13] incorporated local self-attention and convolutional layers, improving the extraction of fine-grained local motion features. Beyond attention mechanisms, RIO [25] introduced rotationally equivariant networks for self-supervised pose estimation, while EqNIO [26] learned rotation-invariant displacement priors through canonical frame mapping. [27]proposed using Lie Events as an alternative to conventional raw IMU data as inputs to the Neural Differential Pose (NDP) model. [28]modeled the dominance of different signals under various motion patterns and generated virtual training samples based on this modeling, thereby alleviating the problem of insufficient labeled data. These advances highlight the rapid pace of innovations in neural network technologies for pedestrian inertial odometry. However, applying deep learning–based inertial odometry to quadruped robots has been far less explored, particularly under the constraint of relying solely on IMU measurements.
The Extended Kalman Filter (EKF) is a widely used extension of the classical Kalman Filter for nonlinear systems, where local linearization is performed through a first-order Taylor expansion. Early works applied EKF to quadruped robot state estimation by fusing IMU and kinematic data. For example, [29] fused IMU and hexapod kinematics but assumed that three legs remained in contact with the ground, limiting generality. Later, [30] combined IMU, joint encoders, and foot-contact sensors within an EKF, while [31] proposed mounting additional IMUs on the legs and fusing them with body-mounted IMU and encoder data, enabling more accurate estimation of both body and foot states. While effective, these methods depend on multiple heterogeneous sensors, which increase hardware and computational complexity.
@@
TLIO [10] bridged learning and filtering by feeding neural network–predicted displacement and uncertainty into an EKF, significantly improving the robustness of IMU-only odometry. This paradigm has since been extended to pedestrian and UAV navigation [16, 18, 17], demonstrating the benefits of combining data-driven models with probabilistic filtering. In the quadruped domain, [15] integrated neural network–based state predictions with joint encoder measurements within an EKF for improved estimation. However, the approach relied heavily on encoder data and introduced little innovation in network design, which reduces generality and practical deployment.
@@
Figure 2: Overall framework of the proposed X-IONet framework. The raw inertial data are rotated using the attitude estimated by EKF and then fed into a rule-based expert selection network to identify the motion platform. The data are routed to a displacement prediction network to regress displacement and uncertainty. The predicted results are refined through the EKF to achieve accurate and robust cross-platform inertial odometry.
@@ III-A Rule-based Expert Selection Network
All convolutional layers use Kaiming normal initialization to ensure stable signal propagation and prevent gradient vanishing or explosion. After the final convolutional layer, adaptive average pooling produces a fixed-length feature vector, which is then fed into a fully connected layer to generate the prediction logits for two classes.
Expert selection module. A lightweight selection module is implemented based on the output of the classification network. The network generates prediction logits that are converted into probabilities via a softmax function, yielding a platform category and its confidence score—class 1 for human and class 0 for quadruped robot. A simple if–else rule then directs each IMU sequence to its corresponding pre-trained displacement estimation expert network. This design is computationally efficient and provides an effective means for generalizing inertial odometry across heterogeneous motion platforms.
@@ III-B Displacement Prediction Network
Inertial odometry regresses motion states from sequential IMU measurements. Since accelerations and angular velocities along each axis are correlated yet heterogeneous, capturing both temporal and inter-axis dependencies is essential for accurate motion estimation. Inspired by Crossformer [32], we design a dual-stage attention enhanced motion feature encoder–decoder architecture to jointly learn these dependencies.
Embedding layer. Given an inertial sequence $\boldsymbol{\mathcal{X}}\in\mathbb{R}^{6\times L}$ of length $L$, we first partition it into equal sub-segments of length $L_{seg}$ for each dimension. The $i$-th segment of the $j$-th dimension, $\boldsymbol{\mathcal{X}}^{seg}_{i,j}$, is linearly projected and augmented with a learnable positional embedding:
@@
where $\boldsymbol{H}\in\mathbb{R}^{d_{model}\times L_{seg}}$ is the projection matrix and $\boldsymbol{E}^{pos}_{i,j}\in\mathbb{R}^{d_{model}}$ is the positional encoding.
Motion Feature Encoder with Dual-Stage Attention. The resulting 2D feature tensor $\boldsymbol{Y}\in\mathbb{R}^{N\times D\times d_{model}}$ (with $N$ segments and $D$ dimensions) is processed through two successive attention stages:
  • Temporal Self-Attention: Multi-head self-attention is applied independently to each dimension to capture long-range temporal dependencies:
@@
  • Dimensional Self-Attention: Multi-head self-attention is then performed across dimensions to model inter-dimensional relationships. A lightweight router [32] selects the most relevant dependencies, reducing computational costs.
@@
The encoder consists of $N$ layers, each merging every two adjacent temporal features (except for the first layer) to obtain coarser temporal representations. Each layer applies the dual-stage attention module to extract multi-scale temporal and dimensional features, enabling hierarchical motion understanding. After encoding, we obtain multi-resolution representations $\boldsymbol{Y}^{enc,0}$, $\boldsymbol{Y}^{enc,1}$, …, $\boldsymbol{Y}^{enc,N}$($\boldsymbol{Y}^{enc,0}=\boldsymbol{H}$), where higher layers capture long-term dynamics and lower layers retain fine-grained motion details.
@@
where the $\delta=0.005$, $\boldsymbol{\hat{d}}_{i}$ and $\boldsymbol{d}_{i}$ denote the predicted and ground-truth displacements and $\hat{\Sigma}_{i}$ is $3\times 3$ covariance matrix.
This formulation provides robustness to outliers via the Huber loss, while the Gaussian likelihood term enables uncertainty-aware weighting of predictions. The Huber-Gaussian loss approximates the Mean Squared Error (MSE) for small errors, ensuring fitting accuracy, and transitions to a linear penalty for large errors, effectively mitigating the impact of outliers. This characteristic contributes to more stable training and better generalization in cross-platform scenarios. Compared to the loss function used in [10], the Huber-Gaussian loss simultaneously optimizes both the size of prediction errors and the accuracy of uncertainty estimation, leading to more stable statistical consistency of the uncertainty (Section IV-D).
@@ III-C1 State Model
The filter state vector is defined as: $\boldsymbol{S}=(\boldsymbol{\lambda_{1}},\dots,\boldsymbol{\lambda_{n}},\boldsymbol{x})$. The historical state vector $\boldsymbol{\lambda}_{i}(i=1,\ldots,n)$ contains pose information from $n$ past moments, and $\boldsymbol{x}=\left(\mathrm{{}_{i}^{w}R},\mathrm{{}^{w}}\boldsymbol{v},\mathrm{{}^{w}}\boldsymbol{p},b_{g},b_{a}\right)$ represents the current state. Where $\mathrm{{}_{i}^{w}}\mathrm{R}\in SO(3)$ is the rotation matrix from the inertial frame to the world frame; $\boldsymbol{\mathrm{{}^{w}}v}$ and $\boldsymbol{\mathrm{{}^{w}}p}$ are the velocity and position in the world frame; $b_{g}$ and $b_{a}$ are the biases of the gyroscope and accelerometer. The EKF uses an error state representation:
@@
TABLE I: Results comparison of different models on Go2 dataset.
Model Easy 01 Easy 02 Medium 01 Medium 02 Hard 01 Hard 02 Average
ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE
RoNIN-ResNet 2.43 1.57 1.98 1.82 4.23 2.13 3.89 1.83 2.07 1.35 7.08 3.00 3.62 1.95
RoNIN-TCN 4.7 3.96 2.39 2.22 2.98 2.11 1.64 1.45 4.88 3.03 5.05 1.98 3.61 2.46
RONIN-LSTM 2.67 2.46 1.58 1.26 3.32 1.44 3.36 1.41 4.85 2.69 8.08 3.23 3.98 2.08
TLIO 3.34 1.92 2.13 1.72 1.27 0.82 1.62 0.76 1.27 1.18 3.66 1.76 2.22 1.37
IMUNet 3.93 2.57 1.45 1.18 0.90 0.84 0.57 0.66 5.39 2.96 6.75 2.40 3.17 1.77
MobileNet 2.67 2.32 2.58 1.46 4.28 2.49 4.20 2.27 1.59 1.46 6.00 2.02 3.56 2.07
MnasNet 7.53 3.71 1.19 1.09 2.38 1.26 0.98 1.08 4.61 2.37 3.97 2.07 3.45 1.93
EfficientNet 3.46 2.03 0.90 0.94 1.41 1.08 1.77 1.13 3.33 2.09 2.23 1.32 2.18 1.43
CTIN 3.42 2.04 3.22 3.00 4.26 2.21 2.49 1.00 1.78 1.04 6.07 2.85 3.54 2.03
Ours 0.85 0.84 0.84 0.91 0.82 0.45 0.50 0.29 1.36 1.05 1.47 1.16 1.03 0.84

### v2

# X-IONet: Cross-Platform Inertial Odometry Network for Pedestrian and Legged Robot

@@ Abstract
Learning-based inertial odometry has achieved remarkable progress in pedestrian navigation. However, extending these methods to quadruped robots remains challenging due to their distinct and highly dynamic motion patterns. Models that perform well on pedestrian data often experience severe degradation when deployed on legged platforms. To tackle this challenge, we introduce X-IONet, a cross-platform inertial odometry framework that operates solely using a single Inertial Measurement Unit (IMU). X-IONet incorporates a rule-based expert selection module to classify motion platforms and route IMU sequences to platform-specific expert networks. The displacement prediction network features a dual-stage attention architecture that jointly models long-range temporal dependencies and inter-axis correlations, enabling accurate motion representation. It outputs both displacement and associated uncertainty, which are further fused through an Extended Kalman Filter (EKF) for robust state estimation. Extensive experiments on the public RoNIN pedestrian dataset, the GrandTour quadruped dataset, and a self-collected Go2 quadruped dataset demonstrate that X-IONet achieves state-of-the-art performance, reducing ATE and RTE by 14.3% and 11.4% on RoNIN, 11.8% and 9.7% on GrandTour, and 52.8% and 41.3% on Go2. These results highlight X-IONet’s effectiveness for accurate and robust inertial navigation across both human and legged robot platforms.
@@ I Introduction
Inertial navigation estimates ego-motion and position solely from Inertial Measurement Unit (IMU) data. Unlike visual–inertial odometry [10, 27, 3], lidar–inertial odometry [31, 20, 8], or GNSS-based localization [17, 25], inertial navigation is fully self-contained and does not rely on external signals [7]. This independence enables robust performance in visually degraded environments and GNSS-denied settings. However, unavoidable sensor noise causes significant drift during long-term pure-inertial integration, making accurate standalone inertial navigation a long-standing challenge.
@@
Recently, deep learning has shown great potential to enhance inertial odometry. Learning-based methods leverage neural architectures to directly extract motion patterns from IMU signals, effectively constraining dead-reckoning drift and improving state estimation accuracy. These approaches have been applied to pedestrian motion [6, 23, 34, 29, 26, 15, 30], legged robots [4], and unmanned aerial vehicles (UAVs) [28, 9, 35, 32].
@@
Despite promising progress, most existing methods are tailored to a single platform, leaving cross-platform inertial odometry largely unsolved. Although pedestrian and quadruped locomotion share certain characteristics—such as periodicity and gravity-related priors—quadruped motion is considerably more dynamic and diverse. As shown in Fig. 1, quadrupeds perform rapid accelerations and decelerations, lateral and backward movements, and highly non-periodic maneuvers, producing complex inertial signatures absent in human motion. Moreover, inertial motion cues are inherently multi-scale and multi-dimensional; existing architectures struggle to extract and fuse such features effectively, resulting in inaccurate displacement predictions. These challenges call for new neural models that capture both long-range temporal dependencies and cross-axis correlations in IMU data.
Figure 1: Overview of the proposed X-IONet model. IMU sequences are first processed by an expert selection network, which identifies the motion platform and routes the input to the corresponding platform-specific motion network. Each motion expert network subsequently predicts the motion trajectory for its designated platform.
@@
To address these challenges, we propose X-IONet (Cross-Platform Inertial Odometry Network), a novel and generalizable deep learning framework for inertial odometry estimation across both human and quadruped robotic platform. As shown in Fig. 1, X-IONet incorporates a rule-based expert selection module that routes IMU data to specialized expert networks based on platform characteristics. Each expert network adopts a dual-stage attention–based displacement prediction model, capturing long-range temporal dependencies via temporal self-attention and inter-axis relationships via dimensional self-attention. The network predicts motion displacement and corresponding uncertainty, and integrates these estimates into an Extended Kalman Filter (EKF) for robust motion state estimation. Extensive experiments on the RoNIN[15] pedestrian dataset, the GrandTour[13] quadruped dataset, and our self-collected Go2 quadruped dataset demonstrate that X-IONet achieves state-of-the-art accuracy, generalizes effectively across platforms, and maintains robustness under distinct motion patterns.
@@
  • We perform comprehensive evaluations on public pedestrian and quadruped datasets, as well as a self-collected quadruped dataset, showing that X-IONet generalizes across distinct motion patterns and achieves state-of-the-art inertial navigation performance.
@@ II RELATED WORK
Recent advances in deep learning–based inertial odometry focus on modeling raw IMU signals for accurate motion estimation. IONet [6] first demonstrated the feasibility of learning motion patterns directly from raw inertial data using recurrent neural networks. RoNIN [15] further advanced this direction by proposing three end-to-end frameworks based on Residual Networks (ResNet) [14], Long Short-Term Memory (LSTM) networks [16], and Temporal Convolutional Networks (TCN) [1], and also released a large-scale dataset for pedestrian inertial navigation. With the success of Transformer architectures in natural language processing [11] and computer vision [12], CTIN [29] introduced a Transformer-based model for inertial data, enabling global motion pattern modeling through dynamic attention computation. Building upon this idea, iMoT [26] incorporated local self-attention and convolutional layers, improving the extraction of fine-grained local motion features. Beyond attention mechanisms, RIO [5] introduced rotationally equivariant networks for self-supervised pose estimation, while EqNIO [19] learned rotation-invariant displacement priors through canonical frame mapping. Jayanth et al.[18] proposed using Lie Events as an alternative to conventional raw IMU data as inputs to the Neural Differential Pose (NDP) model. Li et al.[21] modeled the dominance of different signals under various motion patterns and generated virtual training samples based on this modeling, thereby alleviating the problem of insufficient labeled data. These efforts highlight rapid progress in pedestrian inertial odometry, but deep learning for quadruped robots using only IMU remains underexplored.
@@
The Extended Kalman Filter (EKF) is widely used for nonlinear state estimation. Early works applied EKF to quadruped robot state estimation by fusing IMU and kinematic data. For example, Lin et al.[22] fused IMU and hexapod kinematics but assumed that three legs remained in contact with the ground, limiting generality. Later, Bloesch et al.[2] combined IMU, joint encoders, and foot-contact sensors within an EKF, while Yang et al.[33] proposed mounting additional IMUs on the legs and fusing them with body-mounted IMU and encoder data, enabling more accurate estimation of both body and foot states. While effective, these methods depend on multiple heterogeneous sensors, which increase hardware and computational complexity. TLIO [23] bridged learning and filtering by feeding neural network–predicted displacement and uncertainty into an EKF, significantly improving the robustness of IMU-only odometry. This paradigm has since been extended to pedestrian and UAV navigation [28, 35, 9, 32], demonstrating the benefits of combining data-driven models with probabilistic filtering. In the quadruped domain, Buchanan et al.[4] integrated neural network–based state predictions with joint encoder measurements within an EKF for improved estimation. However, the approach relied heavily on encoder data and introduced little innovation in network design, which reduces generality and practical deployment.
@@
Figure 2: The framework of the proposed X-IONet model. The raw inertial data are rotated using the attitude estimated by EKF, so that the input is represented in a gravity-aligned local frame. The transformed IMU sequence is then fed into a rule-based expert selection network to identify the motion platform. The data are routed to a displacement prediction network to regress displacement and uncertainty. The predicted results are refined through the EKF to achieve accurate and robust cross-platform inertial odometry.
@@ III-A Rule-based Expert Selection Network
All convolutional layers use Kaiming normal initialization to ensure stable signal propagation and prevent gradient vanishing or explosion. After the final convolutional layer, adaptive average pooling produces a fixed-length feature vector, which is then fed into a fully connected layer to generate the prediction logits for two classes. The detailed architecture of the 1-D CNN classifier is illustrated in Fig. 3.
Figure 3: Architecture of the 1-D CNN classifier used in the rule-based expert selection network.
Expert selection module. A lightweight selection module is implemented based on the output of the classification network. The network generates prediction logits that are converted into probabilities via a softmax function, yielding a platform category and its confidence score. A simple if–else rule then directs each IMU sequence to its corresponding pre-trained displacement estimation expert network. The classifier is trained in advance using labeled inertial data from the platform categories to be recognized and is optimized with the standard Cross-Entropy loss[24]. Extending the framework to new platforms requires adding representative training data, retraining the selector, and learning a corresponding expert. This design is computationally efficient and provides an effective means for generalizing inertial odometry across heterogeneous motion platforms.
@@ III-B Displacement Prediction Network
Inertial odometry regresses motion states from sequential IMU measurements. From a physical perspective, accelerometer readings contain gravity-related components and motion-induced specific force, while gyroscope readings describe the rotational dynamics of the carrier. These quantities are not independent: roll and pitch affect the projection of gravity onto each accelerometer axis, and are in turn correlated with angular velocities through the platform motion. Since accelerations and angular velocities along each axis are correlated yet heterogeneous, capturing both temporal and inter-axis dependencies is essential for accurate motion estimation. Inspired by Crossformer [36], we design a dual-stage attention enhanced motion feature encoder–decoder architecture to jointly learn these dependencies.
Embedding layer. Given an inertial sequence $\boldsymbol{\mathcal{X}}\in\mathbb{R}^{6\times L}$ of length $L=200$, we first partition it into equal sub-segments of length $L_{seg}=10$ for each dimension. The $i$-th segment of the $j$-th dimension, $\boldsymbol{\mathcal{X}}^{seg}_{i,j}$, is linearly projected and augmented with a learnable positional embedding:
@@
where $\boldsymbol{H}\in\mathbb{R}^{d_{model}\times L_{seg}}$ is the projection matrix, $\boldsymbol{E}^{pos}_{i,j}\in\mathbb{R}^{d_{model}}$ is the positional encoding, and $d_{model}=256$ denotes the embedding size.
Motion Feature Encoder with Dual-Stage Attention. The resulting 2D feature tensor $\boldsymbol{Y}\in\mathbb{R}^{N\times D\times d_{model}}$ (with $N=20$ segments and $D=6$ dimensions) is processed through two successive attention stages:
  • Temporal Self-Attention: To capture long-range temporal dependencies within each sensor dimension, the feature slice corresponding to the $d$-th dimension $Y_{:,d}\in\mathbb{R}^{N\times d_{model}}$ is processed by a temporal self-attention mechanism:
@@
    where
    $Q=Y_{:,d}W_{Q},K=Y_{:,d}W_{K},V=Y_{:,d}W_{V}$ (4)
    $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d_{k}\times d_{model}}$ are learnable projection matrices. $d_{k}=\frac{d_{model}}{h}$ denotes the dimension of each attention head, with $h$ being the number of heads. This operation models temporal correlations within each dimension independently.
  • Dimensional Self-Attention: After temporal encoding, multi-head self-attention is further performed across dimensions to model inter-dimensional dependencies among sensor channels. For each temporal position $n$, the corresponding feature vector $Y_{n,:}\in\mathbb{R}^{D\times d_{model}}$ is processed by a dimensional self-attention mechanism. A lightweight router [36] is introduced to selectively aggregate the most relevant inter-dimensional dependencies, reducing computational overhead while preserving representational richness.
@@
The encoder consists of $N$ layers, each merging every two adjacent temporal features (except for the first layer) to obtain coarser temporal representations. Each layer applies the dual-stage attention module to extract multi-scale temporal and dimensional features, enabling hierarchical motion understanding. After encoding, we obtain multi-resolution representations $\boldsymbol{Y}^{enc,0}$, $\boldsymbol{Y}^{enc,1}$, …, $\boldsymbol{Y}^{enc,N}$, where higher layers capture long-term dynamics and lower layers retain fine-grained motion details.
@@
where $\delta=0.005$, $\hat{d}_{i}$ and $d_{i}$ denote the predicted and ground-truth displacements, respectively, and $\hat{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ denotes the covariance matrix predicted by the network for the $i$-th sample.
This formulation provides robustness to outliers via the Huber loss, while the Gaussian likelihood term enables uncertainty-aware weighting of predictions. The Huber-Gaussian loss approximates the Mean Squared Error (MSE) for small errors, ensuring fitting accuracy, and transitions to a linear penalty for large errors, effectively mitigating the impact of outliers. This characteristic contributes to more stable training and better generalization in cross-platform scenarios. Compared to the loss function used in [23], the Huber-Gaussian loss simultaneously optimizes both the size of prediction errors and the accuracy of uncertainty estimation, leading to more stable statistical consistency of the uncertainty (Section IV-E).
@@ III-C1 State Model
The filter state vector is defined as: $\boldsymbol{S}=(\boldsymbol{\lambda_{1}},\dots,\boldsymbol{\lambda_{n}},\boldsymbol{x})$. The historical state vector $\boldsymbol{\lambda}_{i}(i=1,\ldots,n)$ contains pose information from $n$ past moments, and $\boldsymbol{x}=\left(\mathrm{{}_{i}^{w}R},\mathrm{{}^{w}}\boldsymbol{v},\mathrm{{}^{w}}\boldsymbol{p},b_{g},b_{a}\right)$ represents the current state. Where $\mathrm{{}_{i}^{w}}\mathrm{R}\in SO(3)$ is the rotation matrix from the IMU frame to the world frame; $\boldsymbol{\mathrm{{}^{w}}v}$ and $\boldsymbol{\mathrm{{}^{w}}p}$ are the velocity and position in the world frame; $b_{g}$ and $b_{a}$ are the biases of the gyroscope and accelerometer. The EKF uses an error state representation: